OpenStudy (anonymous):

Solve. show all work u2 − 2u − 2 = 0

OpenStudy (dumbcow):

Solve using completing the square move 2 to other side u^2 -2u = 2 take half of middle coefficient and square it, add to both sides u^2 -2u +1 = 2+1 Factor left side (u-1)^2 = 3 Square root of both sides u-1 = +-sqrt(3) u = 1 +- sqrt(3)

OpenStudy (dumbcow):

you could also use quadratic formula :)

OpenStudy (anonymous):

o ok thank you for telling me